I am having a hard time with this problem. I can't remember where to begin. The problem is as follows: Graph the linear equation... -x=5y-2 Where do I begin? Thank you in advance for your help.
Answer by lwsshak3(6505)

I am having a hard time with this problem. I can't remember where to begin. The problem is as follows: Graph the linear equation... -x=5y-2 Where do I begin?
**
Change equation into standard form for a straight line: y=mx+b, m=slope, b=y-intercept
-x=5y-2
5y=-x+2
y=-x/5+2/5
This gives you the slope of 1/5 and y-intercept of 2/5
To graph line:
Starting from the y-intercept(0,2/5), move up one unit and 5 units to the right (rise/run) to establish the 2nd point. Drawing a line thru these two points gives you a graph of the given equation.